OpenStudy (anonymous):

What is the fifth term of the geometric sequence? a1 = 120, a2 = 36, a3 = 10.8, a6 = 0.2916

OpenStudy (anonymous):

a2/a1=a3/a2=a5/a6

OpenStudy (anonymous):

@mathsmind Thanks !

OpenStudy (anonymous):

u r welcome!

OpenStudy (anonymous):

wait i don't understand can someone explain it?

OpenStudy (anonymous):

.972 is the answer